Velocity of a particle is expressed as
Vx=80-10t²m/s
(a)=find average acceleration
(b)=instantaneous acceleration
1
Expert's answer
2020-03-17T09:22:49-0400

(a) The average acceleration


"a_x=\frac{v_2-v_1}{t_2-t_1}=\frac{(80-10t_2^2)-(80-10t_1^2)}{t_2-t_1}=-10(t_2+t_1)."

(b) The instantaneous acceleration


"a_x(t)=v_x'(t)=(80-10t^2)'=-20t."